### v2

# Haifa, 3200002, Israel 2]Institut für Physik und Astronomie, Universität Potsdam Karl-Liebknecht-Str. 24/25, 14476 Potsdam-Golm, Germany

Abstract
Hypervelocity white dwarfs (HVWDs) are stellar remnants moving at speeds exceeding the Milky Way’s escape velocity. The origins of the fastest HVWDs are enigmatic, with proposed formation scenarios facing challenges explaining both their extreme velocities and observed properties. Here we report a three-dimensional hydrodynamic simulation of a merger between two hybrid Helium-Carbon-Oxygen white dwarfs (HeCO WDs with masses of 0.69 and 0.62 M⊙). We find that the merger leads to a partial disruption of the secondary WD, coupled with a double-detonation explosion of the primary WD. This launches the remnant core of the secondary WD at a speed of km s-1, consistent with observed HVWDs. The low mass of the ejected remnant and its heating from the primary WD’s ejecta explain the observed luminosities and temperatures of hot HVWDs, which are otherwise difficult to reconcile with previous models (such as the D6). This discovery establishes a new formation channel for HVWDs and points to a previously unrecognized pathway for producing peculiar Type Ia supernovae and faint explosive transients.
Hypervelocity white dwarfs (HVWDs [1, 2, 3, 4]) are enigmatic objects, hurtling through our galaxy at speeds exceeding the Milky Way’s escape velocity[1, 2, 5, 3, 6, 4]. Their existence challenges our understanding of stellar evolution and the dynamics of close binary systems. While most white dwarfs exhibit velocities typical of the Galactic disk or halo, HVWDs possess extreme speeds, implying a violent ejection mechanism. The advent of high-precision astrometry from the Gaia mission [7, 8] has led to the identification of seven HVWD candidates with velocities exceeding 1000 km/s [3, 4], further deepening the mystery surrounding their origin. Of these, two of the first identified HVWDs (D6-2 and D6-3) have very low measured radial velocities (RVs), possibly inconsistent with the far higher measured tangential velocities, casting doubt on their identification as HVWDs and/or their measured velocities [3, 9, 10], while the other five have large radial velocities consistent with typical km s-1 (see Supplementary Table 1). As they had more time to be excited by galactic perturbations, old WDs might also have non-negligible peculiar velocities of up to a few hundred km s-1, suggesting a larger uncertainty in ejection velocity. In particular, in-going (towards the direction of the Galactic disk) HVWDs could not have originated from the disk, suggesting large ages and peculiar velocities for their progenitors at the time of ejection. D6-2 and J0927 are in-going HVWDs, which might suggest that the inferred extreme ejection velocity of J0927 (the highest velocity inferred; [4]) could be lower, more in par with other HVWDs, and potentially closer to km s-1, but conversely, could also be higher.
Lower velocity HVWDs with velocities of up to hundreds of km s-1 were suggested to be produced from partial deflagration supernovae (SNe) of WDs and indeed, such objects were later identified; see [1, 2, 4, 11, 12] and references therein. [13] suggested the possibility of ejecting high-velocity evolved stars with velocities up to 100 km s-1 in the context of the single degenerate scenario for type Ia SNe, i.e., a Blaauw-kick scenario for companions of exploding WDs[14]). A leading hypothesis for the formation of bona-fide HVWDs with km s-1 follows a similar direction of a Blaauw-kick following a SN in a binary system, but for a different model for type Ia SNe, in which the ejected companion is a WD. It involves the ”dynamically driven double-degenerate double-detonation” (D6) scenario [15, 16, 17, 3] (based on earlier double detonation scenarios not involving two WDs [18, 19]). In this model, a close binary system consisting of two white dwarfs (one typically massive ) undergoes a gravitational wave inspiral. As the primary white dwarf accretes a small amount of Helium from its companion, a Helium detonation is triggered in the accreted and any low-mass pre-existing Helium shell, leading to a second, carbon detonation in the core and a Type Ia SN explosion. The companion white dwarf, no longer gravitationally bound, is flung outward at high speed. As we discuss in detail below, while we also consider the case of a merger of two WDs, we focus on low-mass WDs, which show a different behavior, and our findings differ significantly from the D6 model in this context.
Nevertheless, the D6 scenario faces difficulties in explaining both the extreme HVWD velocities ( 2000 km s-1) and the observed properties of these objects[20]. Recent studies [11] have shown that detonation of a massive WD primary scenario as envisioned in the D6 scenario does not produce the high temperatures and inflated radii on observable timescales in some HVWDs, particularly for those with the highest velocities, for which the D6 scenario requires high mass HVWDs ( M⊙[21, 20]). Moreover, the predicted rate of HVWD formation from D6 events, if this were the primary channel for the origin of Type Ia supernovae, appears to be significantly higher than the observed rate[3, 9]. Additionally, HVWDs have not been found in close proximity to secure Type Ia supernova remnants[9, 22], although a possible association has been noted for one case[3]. More recent studies suggested that companions might also explode[23, 24, 25, 26], which could potentially resolve the rates and supernova remnant issue, without significantly affecting the SN observations [25, 27, 26]. However, in this case, no HVWD would form at all, requiring an alternative origin, like the one suggested here.
Here, we present a novel scenario for the origin of HVWDs based on a three-dimensional hydrodynamical simulation of a merger between two hybrid Helium-Carbon-Oxygen white dwarfs (HeCO WDs). These WDs are likely the remnants of previously interacting stars that underwent a common-envelope (CE) phase, resulting in a thick He shell surrounding a degenerate CO core (see [28, 29] and references therein). Our simulation reveals a previously unexplored pathway for HVWD formation, distinct from the standard D6 model. We find that the merger of two HeCO WDs (0.68 and 0.62 M⊙) can lead to a partial tidal disruption of the lower-mass white dwarf (i.e. the secondary or companion), coupled with a double-detonation explosion of the primary. This process launches the remnant low-mass hot core of the disrupted white dwarf at hypervelocity speeds, which can remain inflated and hot for long timescales, consistent with observations. Thereby, our findings provide a compelling explanation for the origin of the fastest HVWDs and offer new insights into the diversity of explosive stellar phenomena.
We model the merger using 3D hydrodynamical simulation using the Arepo code[30] (see Methods). Figure 1 illustrates the key stages of the merger and explosion. A full movie showing the merger can be found in the SI (Supplementary Video 1), as well as a figure with more detailed snapshots (Sect. 2.2). As the secondary WD approaches the primary, it is tidally deformed and begins to transfer mass. This mass transfer eventually triggers a helium detonation on the surface of the primary WD. In the D6 scenario the high temperature and density conditions on the massive primary lead to early Helium detonation, occurring while the secondary is still intact, and after minimal accretion from the companion. However, in the low-mass HeCO cases shown here, a Helium detonati on occurs only after the secondary has already begun to be significantly disrupted. In particular, a significant mass of Helium has already been transferred from the secondary to the primary, allowing for stronger, more energetic Helium detonation. The Helium detonation propagates around the primary, converging on the opposite side and driving a shock wave into the core. This shock wave initiates a secondary detonation in the Carbon-Oxygen core, leading to a supernova explosion, leaving the secondary unbound and ejected at a higher velocity. The very close approach and partial disruption allow for higher orbital (and eventually ejection) velocities and stronger heating of the companion. We note that since we followed the evolution for 200s, and the relatively low mass and density remnant has not been detonated, we don’t expect it to ignite, as it further cools down in the long term, nor do we find evidence for this in our models of the long-term evolution discussed below.
Although the secondary WD is partially disrupted during the merger (stripped of M⊙, and transferred M⊙ to the primary), its core remains intact, when the Carbon-Oxygen detonation ensues. This core, tidally and debris-impact-heated and now unbound, is ejected following the supernova explosion at a velocity of 2061 km s-1. The properties of the ejected remnant are summarized in Table 1. The remnant has a mass of 0.492 M⊙, including a small amount of Helium (0.003 M⊙) and traces of heavier elements synthesized during the explosion. Due to the strong tidal interaction during the merger, the ejected HVWD is expected to be rapidly rotating. The initial rotation velocity should be comparable to the ejection velocity. However, as the remnant expands and cools, its rotation velocity decreases due to angular momentum conservation. At the final snapshot of our simulation, the HVWD has a total angular momentum of g cm2 s-1; during its long-term evolution, the WD expands and its rotational velocity is expected to be significantly lower. For the currently inferred radius of R⊙ for some of the HVWDs, we would expect a rotation velocity of km s-1.
| Kick Velocity | Angular Momentum | ||||
| [km s-1] | [g cm2 s-1] | ||||
| 0.49 | 2061 |
The supernova explosion resulting from the double detonation releases erg of nuclear energy (Table 2). The explosion ejects a total of 0.82 M⊙ of material, including 0.1 M⊙ of iron-group elements, of which 0.072 M⊙ is radioactive 56Ni, and 0.42 M⊙ of intermediate-mass elements. Interestingly, the Helium shell burning and detonation prior to the CO detonation already produces a few 0.01 M⊙ of both radioactive 56Ni, as well as faster decaying radioactive elements 48Cr and 52Fe which could affect the early light-curve, and typically not expected in normal type Ia SNe (see the radioactive elements production evolution in time in Figure 2 and the final composition in Supplementary Table 2). This finding agrees with previous He-detonation studies [31, 32, 33] with similar element production in their corresponding cases, even though our heavier 0.59 M⊙ CO core has less He before the first detonation (about 0.13 M⊙), (see Supplementary Table 3 for the element produced in the first (He) detonation,). Comparing the total element production with previous works, we find that despite the differences in total mass and He shells, as well as differences in the numerical codes, our results also lie in the same regime as in the more similar double-detonation cases of Sim et al. 2010[32] (see Supplementary Table 2). We note that it is still unclear what the conditions result in double detonation vs. a single He-shell detonation. Generally, double-detonation is expected to occur when both the He-shell and the CO core are denser, typically corresponding to higher WD mass, and when the He detonation is stronger, and produces a stronger shock into the CO core.
| Total Nuclear Energy | SN CoM Velocity | |||||
| erg] | [ | [km s-1] | ||||
| 0.68 | 0.14 | 0.82 | 0.1 | 0.42 | 1187 |
Table 2 summarizes the SN explosion properties. Given the low 56Ni produced, the low ejecta velocity (median 7000-8000 km s-1),, and the relatively small amount of ejecta, the SN is likely to be faint and peculiar in respect to normal type Ia SNe. These properties (low mass, low 56Ni and low ejecta velocity)might be consistent with 91bg-like or 02es-like SNe [34]. The low abundances of radioactive elements may give rise to low photospheric temperatures, and together with the large 44Ti abundance produced in the outer ejecta (see Supplementary Table 3 and Supplementary Table 2) could potentially induce early prominent Ti lines in the spectra, consistent with 91bg-like SNe. The properties of this SN might also resemble some properties of a Ca-rich SN, e.g. large abundances of intermediate elements such as 44Ti and 40Ca (see Supplementary Table 2). However, the 56Ni mass produced, and the total energetics are a few times higher than typical Ca-rich SNe. We do note that all of these sub-types of SNe are preferentially observed in old environments in early-type galaxies [35, 34]. In addition, since the primary also had a large orbital velocity at the time of the explosion, the SN produced should have a large center-of-mass velocity. This could potentially be identified as a shift in the velocity measurement of nebular spectra taken long after the explosion of such a peculiar, faint SN. The supernova remnants (SNR) of such SNe, if identified, should also have such a shift. In particular, if paired with an identified HVWD, it should have a velocity in a counter direction to that of the HVWD. We find the SNR center of mass velocity in our simulation to be 1187 km s-1.
Following the simulation, we record the remnant properties (densities, compositions, temperature) found in the simulation (see Table 1), and then follow its long-term evolution, following the same approach as ref. [11], by mapping the properties into 1D and then using the MESA stellar evolution code[36] (see Methods). Figure 3 shows the long-term evolution of the ejected HVWD in the Hertzsprung-Russell diagram (HR diagram). The remnant is initially hot and luminous, and experiences some super-Eddington mass-loss, but then cools and fades over time as it expands and releases its thermal energy. The low mass of the remnant contributes to its inflated state and prolonged cooling time.
The observed physical properties of several HVWD candidates are also plotted in Figure 3, taken from [4], besides for J0927[37]. The radius and temperature of our simulated remnant align well with the observed values, particularly for the HVWD J0546 and J1332, and to a lesser extent with J0927 (consistent within 2-sigma uncertainty, or 1, if ref. [4] inferred temperature is adapted), but see further discussion on the properties J1332 in the SI, section 2.4. This suggests that our proposed formation scenario can explain the observed properties of the more reliable and highest-velocity HVWDs. That being said, D6-1 and J1235 do show consistent reliable high radial velocities but are not directly explained by our specific simulation. Nevertheless, given the large phase space of HeCO WD mergers, it is possible that more consistent models can be found with more/less massive HeCOs configurations (see also discussion on velocities in the SI, section 2.3).
Our 3D hydrodynamical simulation has revealed a new channel for the formation of HVWDs through the merger of two hybrid HeCO white dwarfs. This scenario differs significantly from the previously proposed D6 model, which involves the detonation of a massive CO white dwarf, leading to a normal type Ia SN, following very little mass transfer, leaving the companion intact. In our simulation, the merger of two lower-mass HeCO WDs leads to a partial tidal disruption of the secondary WD and a double-detonation explosion of the low-mass primary. The close approach of the binary down to the point of partial tidal disruption allows for the larger velocity of the companion, even though the primary is relatively low-mass. Hence, this leads to the ejection of the remnant core of the secondary as an HVWD with a velocity exceeding 2000 km s-1. This leads to a fainter SN than a normal type Ia.
This novel channel offers several advantages in explaining the observed properties of HVWDs. First, it naturally produces the high ejection velocities observed for some HVWDs, which are difficult to achieve in the D6 scenario without invoking massive progenitor WDs, which are rare[21, 20]. Second, the low mass of the ejected remnant and its heating from the primary WD’s ejecta explain the observed luminosities and inflated radii of HVWDs, which can not be reconciled with the D6 model for HVWDs with such velocities.
Our findings have broader implications for our understanding of stellar evolution and explosive transients. They highlight the potential importance of HeCO WDs in producing a variety of thermonuclear events. The merger of HeCO WDs may lead to peculiar Type Ia supernovae with fainter luminosities and potentially distinct observational signatures compared to standard SNe Ia. Further studies are needed to characterize the detailed properties of these explosions and their potential contribution to the observed diversity of SNe Ia.
The discovery of HVWDs was initially hailed as a smoking-gun signature of the D6 scenario and its potential role in producing the majority of SNe Ia. However, our results suggest that HVWDs may originate from a wider range of progenitor systems and explosion mechanisms. Together with the various other challenges to the D6 scenario discussed above (e.g., rates and velocities, no evidence for HVWD inside type Ia SNRs), our results cast further doubt on the D6 scenario as the dominant channel for SNe Ia with ejected HVWDs, and emphasize the need to explore alternative pathways.
The current study is limited by the computational cost of 3D hydrodynamical simulations, which restricts the exploration of a wider parameter space of HeCO WD mergers. Future studies with larger parameter surveys will help to constrain the range of HVWD properties and the rates of these events. Nevertheless, our findings demonstrate the viability of HeCO WD mergers as a source of HVWDs and open up new avenues for research into the explosive fates of binary white dwarf systems (see also [38, 24, 33] for other explosive transients involving HeCO WDs). Moreover, analytic estimates (see SI; section 2.3) suggest that it is likely that the phase space of merging double HeCO WDs could produce HVWDs of a wider range of velocities and masses beyond the specific point derived from the fully modeled case shown here. It could naturally explain (and predict) a wider range of velocities in the range km s-1 expected for HVWDs from the partial HeCO tidal-disruption channel. Indeed, the inferred velocities (within the uncertainties) of all besides D6-2 (which is the least reliable given its very low radial velocity) of the HVWDs reside in this range. In fact, these might explain the preference for this velocity range.
In summary, we have demonstrated that the merger of two HeCO WDs can produce HVWDs with properties consistent with observations, both in terms of ejection velocities, ages, and physical properties. This novel channel provides a compelling explanation for the origin of the fastest HVWDs and sheds new light on the diversity of explosive transients in the Universe.
1 Methods
We investigate the formation of HVWDs through the merger of two hybrid Helium-Carbon-Oxygen (HeCO) white dwarfs using a three-dimensional magnetohydrodynamical simulation. This is the first study to explore this scenario, in particular, in 3D, allowing us to capture the complex dynamics of the merger process and the subsequent ejection of the HVWD.
1.1 HeCO White Dwarfs
HeCO white dwarfs are a class of stellar remnants characterized by a Helium-rich outer layer surrounding a Carbon-Oxygen core. Their formation is linked to binary star interactions, where processes like Roche lobe overflow and common envelope evolution lead to the accumulation of Helium on the white dwarf [28, 39, 24]. These hybrid WDs can have masses and compositions that fall between those of typical Helium WDs and Carbon-Oxygen WDs, making their merger outcomes particularly interesting.
More specifically, we investigate the fate of a double degenerate binary system consisting of a primary Helium-rich HeCO WD, with the total mass of made of a CO core of and a massive Helium shell of (see details in [24]), and a secondary hybrid HeCO WD with total mass of made of a CO core of and a massive Helium shell of (following [40]). See Sec. 2.5 for a discussion on the rate of such interactions. Note that the much higher He shell for this hybrid is attained through a somewhat different evolution than that described in [29], as explored in [24].
1.2 Hydrodynamical Simulation
We model the merger of two HeCO white dwarfs using the moving-mesh hydrodynamics code Arepo [30, 41, 42, 43]. This code provides a self-gravitating, adaptive mesh framework for simulating astrophysical phenomena, incorporating a Helmholtz equation of state for degenerate matter and a 55-isotopes nuclear reaction network to capture the thermonuclear processes involved in the merger. A detailed description of WD merger simulations in Arepo can be found in [24], and our models closely follow similar approaches and modeling. We run our simulation with a cell mass resolution of g, a minimum cell volume of cm3, cells, in a total box size of cm to incorporate the explosion and the expansion of the remnant.
1.2.1 Initial Conditions
The initial conditions for our simulation are derived from 1D stellar evolution models. These models provide the density, temperature, and composition profiles for both the primary and secondary HeCO white dwarfs. The specific parameters of our simulated binary system are:
Primary WD: Mass = 0.69 M⊙, CO core = 0.59 M⊙ (0.26 M⊙ 12C and 0.33 16O), He shell = 0.1 M⊙ as modeled in [24] using MESA [36, 44, 45], and mapped to 3D. Secondary WD: Mass = 0.62 M⊙, CO core = 0.58 M⊙ (equal amount of 12C and 16O), He shell = 0.038, M⊙, where the composition is based on MESA models in [39]. This model was generated synthetically in the same method as in [25].
The 3D white dwarf models are initialized in Arepo, relaxed to ensure stability, and then placed in a binary configuration within a large simulation domain. The initial separation of the binary is chosen to be wide enough to minimize initial tidal forces, km, and the corresponding orbital period was s.
1.2.2 Inspiral and Merger
We simulate the inspiral of the binary system, initially accelerating it by artificially removing angular momentum to mimic the effect of gravitational wave emission. This artificial inspiral is halted just before the secondary WD enters the Roche radius of the primary and mass transfer becomes dynamically significant (see [24] for a similar approach and further details). The binary separation at this point was km, with an orbital period of s.
We then follow the fully realistic evolution of the merger. As the secondary WD approaches the primary, it is tidally deformed and disrupted (s after the end of the accelerated inspiral stage). The dynamical mass transfer of Helium from the secondary triggers a Helium detonation on the surface of the primary. This detonation propagates around the primary, leading to a secondary detonation in the Carbon-Oxygen core and a supernova explosion. The partially disrupted secondary WD, still containing its core, is ejected at high velocity by the force of the explosion. These stages are presented in Supplementary Figure 1, while some are also in Figure 1 of the main part. A full movie of the simulation can be found in the SI (Supplementary Video 1).
1.3 Evolution of the Bound Remnant HVWD
To model the long-term evolution of the runaway WD we follow a similar approach as in [11]. We map the output of the simulation ( s after the explosion) to the 1D stellar evolution software MESA. This procedure has been applied before for Athena++ simulations of Helium subdwarfs[21] and more recently of Helium white dwarfs[12], as well as Arepo simulations of massive white dwarfs [11]. Here, we follow the procedure outlined in reference [11]. The 3D profiles of the runaway WD (taken at the last snapshot, which is 10-20 times dynamical timescales after the explosion, allowing sufficient time for the WD to adjust) are averaged to give a 1D profile with temperature, density, and mass fraction information. A 22 isotope network approx_21_plusco56 is used to model the nuclear reactions during this evolution and includes the isotopes 1H, 3,4He, 12C, 14N, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, 44Ti, 48,56Cr, 52,54,56Fe, 56Co, and 56Ni. It therefore includes the most important reactions of nickel decay and capture on Carbon. We relax a white dwarf to our output profiles of composition and entropy, as outlined in [11], and we get rid of almost all of the outer layers where 56Ni dominates the atmosphere to such an extent that the material will become unbound on short timescales (but longer than the hydrodynamical simulation). This cut is chosen again to be the mass coordinate where the energy due to 56Ni decay is twice the gravitational binding energy of the star. This results in a final mass of 0.482 M⊙, with only 0.005 M⊙ cut-off. Once relaxed the white dwarf is allowed to evolve for a maximum time of 100 Myr. We present the results in Figure 3 and Supplementary Figure 2. In addition, as we show in Supplementary Table 2, there is still bound He at the end of our simulation, which is kept there also when following its longer evolution; it might however, keep changing on an even longer timescale. Nevertheless, the existence of a bound He in the remnant will likely lead to He detection in the remnant’s atmosphere.
Data Availability
The MESA input files and inlists for relaxation, heating, and evolution can be accessed at Zenodo [46].
All data underlying this research, which is not in Zenodo (in particular, the many too large, hydrodynamical simulation snapshots), will be available upon reasonable request to the corresponding authors.
Code Availability
Acknowledgments
HBP acknowledges support for this project from the European Union’s Horizon 2020 research and innovation program under grant agreement No 865932-ERC-SNeX. HG acknowledges support for the project from the Council for Higher Education of Israel. A.B. was supported by the Deutsche Forschungsgemeinschaft (DFG) through grant GE2506/18-1 and by the Kavli Summer Program which took place at MPA in Garching in July 2023, and was supported by the Kavli Foundation. We would like to thank Evan Bauer and Robert Fisher for their valuable comments and discussions.
Author contributions statement
HG led the key research, ran the hydrodynamical simulations, and analyzed their results. HBP initiated and supervised the project, suggested the main ideas, analyzed the results, and wrote the main parts of the paper. AB made the long-term WD evolution modeling and analyzed them. RP assisted with running the hydrodynamical simulations. All authors contributed to writing the paper, making the figures, and reviewing the manuscript.
Corresponding authors: Correspondence and requests for materials should be addressed to Hila Glanz or Hagai Perets.
Competing interests statement
The authors declare no competing interests.